Causwell Company began 2018 with 30,000 units of inventory on hand. The cost of each unit was $5.00. During 2018 an additional 50,000 units were purchased at a single unit cost, and 40,000 units remained on hand at the end of 2018 (40,000 units therefore were sold during 2018). Causwell uses a periodic inventory system. Cost of goods sold for 2018, applying the average cost method, is $224,000. The company is interested in determining what cost of goods sold would have been if the FIFO or LIFO methods were used.

Required:
1. Determine the cost of goods sold for 2018 using the FIFO method. [Hint: Determine the cost per unit of 2018 purchases.]
2. Determine the cost of goods sold for 2018 using the LIFO method. (Do not round intermediate calculations.)

CAN YOU PLEASE DO IT QUICKLY AND CORRECTLY? THANK YOU
Answer #1

Calculate purchase cost per unit

Cost of goods sold = Total Cost/Total Unit*Sale unit

224000 = (30000*5+50000*X)/80000*40000

224000 = (150000+50000X)/80000*40000

5.6 = 1.875+0.625X

X = 5.96

1) Cost of goods sold under FIFO = (30000*5+10000*5.96) = 209600

2) Cost of goods sold under LIFO = 40000*5.96 = 238400